5) Suppose that the function P=13+45ln x represents the percentage of inbound e-mail in the U.S. that is considered spam, where x is the number of years after 2000.
Carry all calculations to six decimals on each intermediate step when necessary.
a) Use this model to approximate the percentage of spam in the year 2003.

b) Use this model to approximate the year that the percent of spam will reach 95% provided that law enforcement regarding spammers does not change.

Answer by stanbon(48535)

Suppose that the function P=13+45ln x represents the percentage of inbound e-mail in the U.S. that is considered spam, where x is the number of years after 2000.
Carry all calculations to six decimals on each intermediate step when necessary.
a) Use this model to approximate the percentage of spam in the year 2003.
P=13+45ln x
P(3) = 13 + 45ln3 = 62.437553 %
------------------
b) Use this model to approximate the year that the percent of spam will reach 95% provided that law enforcement regarding spammers does not change.
95 = 13 + 45lnx
82 = 45lnx
lnx = (82/45)
x = e^(82/45)
x = 6.185589
2000 + 6.185589 = 2006.185589
Rounding to the nearest whole number you get the year 2006
==================
Cheers,
Stan H.
---------------